# Asymptotically normal estimators in high-dimensional linear regression

### v1

@@ 4 Proofs
for any $\zeta>0$. As well as the proof of Lemma 2.4, we have
$\,\mathsf{P}\left(n^{-1}\left|\sum_{i=1}^{n}(\tilde{\bm{\theta}}_{n}-\bm{\theta}_{0})_{\hat{T}_{n}}^{\top}\bm{Z}_{i\hat{T}_{n}}\epsilon_{i}\right|>\zeta\right)$
$\leq\,\mathsf{P}\left(\left\{n^{-1}\left|\sum_{i=1}^{n}(\tilde{\bm{\theta}}_{n}-\bm{\theta}_{0})_{T_{\gamma}}^{\top}\bm{Z}_{iT_{\gamma}}\epsilon_{i}\right|>\zeta\right\}\cap\{\hat{T}_{n}=T_{\gamma}\}\right)+\,\mathsf{P}(\hat{T}_{n}\neq T_{\gamma})$
@@
We thus get
$n^{-1}\left|\sum_{i=1}^{n}(\tilde{\bm{\theta}}_{n}-\bm{\theta}_{0})_{T_{\gamma}}^{\top}\bm{Z}_{iT_{\gamma}}\epsilon_{i}\right|\leq\|(\tilde{\bm{\theta}}_{n}-\bm{\theta}_{0})_{T_{\gamma}}\|_{2}\left\|n^{-1}\sum_{i=1}^{n}\bm{Z}_{iT_{\gamma}}\epsilon_{i}\right\|_{2}=o_{p}(1)\cdot O_{p}(1)=o_{p}(1),$
because $n^{1/2}\|\tilde{\bm{\theta}}_{n}-\bm{\theta}_{0}\|_{2}$ is $O_{p}(1)$ and $s_{\gamma}$ is independent of $n$ and $p$. Moreover, Lemma 2.1 implies $\mathsf{P}(\hat{T}_{n}\neq T_{\gamma})=o(1)$. To prove $S_{n3}=o_{p}(1)$, it suffices to show that
$\frac{1}{n}\sum_{i=1}^{n}\left\{(\tilde{\bm{\theta}}_{n}-\bm{\theta}_{0})_{T_{\gamma}}^{\top}\bm{Z}_{iT_{\gamma}}\right\}^{2}=o_{p}(1).$
@@
Next, we show the consistency of $\hat{\Lambda}_{n,1},\ldots,\hat{\Lambda}_{n,r}$. It follows from Weyl’s inequality and Lemma 2.2 that $\max_{j\in[r]}|\hat{\Lambda}_{n,j}-\Lambda_{j}|\leq\|\hat{\bm{\Sigma}}_{nA}-\bm{\Sigma}_{A}\|_{\mathrm{op}}=o_{p}(1)$. ∎
@@
Let $\mathcal{H}_{0}$ hold. Since $\|\bm{R}_{A}\|_{2}^{2}$ has a continuous distribution function, we have $\hat{c}_{\alpha}\to^{p}c_{\alpha}$ by Lemma 3.2. Hence, by Slutsky’s lemma, $(W_{n},\hat{c}_{\alpha})\to^{d}(\|\bm{R}_{A}\|_{2}^{2},c_{\alpha})$. Since $\mathsf{P}(\|\bm{R}_{A}\|_{2}^{2}=c_{\alpha})=0$, we have $1_{\{W_{n}>\hat{c}_{\alpha}\}}\to^{d}1_{\{\|\bm{R}_{A}\|_{2}^{2}>c_{\alpha}\}}$ by the continuous mapping theorem. Finally, it follows from the the Portmanteau theorem that $\mathsf{P}(W_{n}>\hat{c}_{\alpha})\to\,\mathsf{P}(\|\bm{R}_{A}\|_{2}^{2}>c_{\alpha})=\alpha$. ∎
@@ Acknowledgements
The second author was supported by Japan Society for the Promotion of Science KAKENHI Grant Number 25K07133.

### v2

@@ 4 Proofs
for any $\zeta>0$. As well as the proof of Theorem 2.3, the left-hand side admits the bound
$\,\mathsf{P}\left(\left\{n^{-1}\left|\sum_{i=1}^{n}(\tilde{\bm{\theta}}_{n}-\bm{\theta}_{0})_{T_{\gamma}}^{\top}\bm{Z}_{iT_{\gamma}}\epsilon_{i}\right|>\zeta\right\}\cap\{\hat{T}_{n}=T_{\gamma}\}\right)+\,\mathsf{P}(\hat{T}_{n}\neq T_{\gamma})$
@@
We have
$n^{-1}\left|\sum_{i=1}^{n}(\tilde{\bm{\theta}}_{n}-\bm{\theta}_{0})_{T_{\gamma}}^{\top}\bm{Z}_{iT_{\gamma}}\epsilon_{i}\right|\leq\|(\tilde{\bm{\theta}}_{n}-\bm{\theta}_{0})_{T_{\gamma}}\|_{2}\left\|n^{-1}\sum_{i=1}^{n}\bm{Z}_{iT_{\gamma}}\epsilon_{i}\right\|_{2}=o_{p}(1)O_{p}(1)=o_{p}(1),$
because $n^{1/2}\|\tilde{\bm{\theta}}_{n}-\bm{\theta}_{0}\|_{2}=O_{p}(1)$ and $s_{\gamma}$ is independent of $n$ and $p$. Moreover, Lemma 2.1 implies $\mathsf{P}(\hat{T}_{n}\neq T_{\gamma})=o(1)$. To prove $S_{n3}=o_{p}(1)$, it suffices to show that
$n^{-1}\sum_{i=1}^{n}\left\{(\tilde{\bm{\theta}}_{n}-\bm{\theta}_{0})_{T_{\gamma}}^{\top}\bm{Z}_{iT_{\gamma}}\right\}^{2}=o_{p}(1).$
@@
Next, from Weyl’s inequality and Lemma 2.2, we have
$\max_{j\in[r]}|\hat{\Lambda}_{n,j}-\Lambda_{j}|\leq\|\hat{\bm{\Sigma}}_{nA}-\bm{\Sigma}_{A}\|_{\mathrm{op}}\leq\|\hat{\bm{\Sigma}}_{nA}-\bm{\Sigma}_{nA}\|_{\mathrm{op}}+\|\bm{\Sigma}_{nA}-\bm{\Sigma}_{A}\|_{\mathrm{op}}=o_{p}(1),$
where $\bm{\Sigma}_{nA}=\bm{A}_{[q],T_{\gamma}}\bm{J}_{nT_{\gamma},T_{\gamma}}^{-1}\bm{A}_{[q],T_{\gamma}}^{\top}$. ∎
@@
Let $\mathrm{H}_{0}$ hold. Since $\|\bm{R}_{A}\|_{2}^{2}$ has a continuous distribution function, we have $\hat{c}_{\alpha}\to^{p}c_{\alpha}$ by Lemma 3.2. Hence, by Slutsky’s lemma, $(W_{n},\hat{c}_{\alpha})\to^{d}(\|\bm{R}_{A}\|_{2}^{2},c_{\alpha})$. Since $\mathsf{P}(\|\bm{R}_{A}\|_{2}^{2}=c_{\alpha})=0$, we have $\varphi_{n}\to^{d}1_{\{\|\bm{R}_{A}\|_{2}^{2}>c_{\alpha}\}}$ by the continuous mapping theorem. Finally, it follows from the the Portmanteau theorem that $\mathsf{P}(\varphi_{n}=1)\to\,\mathsf{P}(\|\bm{R}_{A}\|_{2}^{2}>c_{\alpha})=\alpha$. ∎
@@ Acknowledgements
The second author was supported by the Japan Society for the Promotion of Science KAKENHI Grant Number 25K07133.